# Robust Channel Estimation for Optical Wireless Communications Using Neural Network

### v1

@@ IV Simulation Results
For the rest of this paper, $M$ is assumed to be 2 which represents a two-path channel for common indoor OWC scenarios. Normalized mean square error (NMSE) is a key performance metric that evaluates the distance between the actual channel and the estimate of channel for each resource element in the slot, which is defined as
$\mathrm{NMSE}(\mathbf{\hat{H}_{Slot}},\mathbf{H})=\frac{\mathbb{E}\left\{\left\Arrowvert\hat{H}_{\mathrm{Slot}}-H\right\Arrowvert_{F}^{2}\right\}}{{\mathbb{E}\{\left\Arrowvert H\right\Arrowvert_{F}^{2}\}}},$ (8)
where $\left\Arrowvert\cdot\right\Arrowvert_{F}^{2}$ is the Frobenius norm. The bit error ratio (BER) is another performance metric which refers to the ratio of bit errors per frame. For the settings of optical wireless channel, the hyper-parameters are given in Table. I. As the CP is shown to have a negligible impact on the electrical signal-to-noise ratio (SNR) requirement for optical systems [18], we consider environment SNRs from 15dB to 30dB in this paper.
TABLE I: System settings for simulations
@@
Modulation schemes 64 QAM
Number of subcarriers $N_{f}$ 324
CP length $L_{CP}$ 7
Pilot spacing $L_{s}$ 5
OFDM symbols in one slot $N_{s}$ 14
DC bias 7dB
Frequency spacing 30kHz
The offline training dataset for the three InterpolateNet networks (LDS, MDS or HDS channels) consists of 100,000 channel samples with the SNR values for the environmental SNR range specified above, 95% for training and 5% for validation. Referring to the simulation settings given in Table. I, both $\mathbf{h_{\mathrm{LDS}}}$ and $\mathbf{h_{\mathrm{HDS}}}$ are assumed to be $\mathbf{h_{\mathrm{LDS}}}$ = 1e-4 * [6.4, 0.21930, 0.09676, 0.06175, 0.04517] and $\mathbf{h_{\mathrm{HDS}}}$ = 1e-4 * [5.5, 0.30126, 0.13441, 0.08609, 0.06310] respectively for this indoor scenario. Moreover, an InterpolateNet trained with HDS condition is implemented for comparison. We use MSE loss for training InterpolateNets and the detailed training hyper-parameters are given in Table. II.
TABLE II: Offline training hyperparameters
Parameters Values
Optimizer Adam
Maximum epoch 100
Initial learning rate (lr) 0.0002
Drop period for lr every 10
Drop factor for lr 0.3
Minibatch size 64
L2 regularization $10^{-9}$
@@ IV-A NMSE performance on randomly generated channels
We evaluate the NMSE performance of the proposed method on randomly generated channels over both SNR and time, to investigate the robustness property for practical implementations. The test channels are generated randomly according to Table. I.
@@
(b) NMSE performance over 90 seconds time with SNR fixed at 20dB.
@@
Figure. 3a shows the NMSE performance over the SNR from 15dB to 30dB and each sample is averaged by 100,000 independent channel realizations. The proposed method is shown to have superior performance than other methods except for MMSE method. At 30dB SNR, the adaptive InterpolateNet achieves a NMSE of 0.5046 which is 0.0020 lower than that of HDS InterpolateNet. The MMSE method has a minimum NMSE value of 0.5001.
For the figure. 3b, the channel will alter among LDS, MDS and HDS for each 10 seconds and each sample is averaged over 100 independent channel realizations. Both the proposed method and the InterpolateNet trained with HDS condition outperform the LS estimate but are worse than MMSE method. The proposed method provides a better NMSE performance with a mean value of 0.5034, which is lower than 0.5052 for the InterpolateNet trained with HDS condition. Compared to the MMSE estimate with NMSE of 0.5012, the proposed solution does not require any actual channel information for online implementation resulting in poorer estimate performance and the NMSE variance of the proposed method is reasonably small over time.
@@ IV-B BER performance on randomly generated channels
We also evaluate the BER performance of each method on randomly generated channels over SNR. The test channels are also generated randomly according to Table. I.
Figure 4: BER performance of LS, MMSE, direct signal detection, InterpolateNet trained with HDS conditions and the proposed method over SNR from 15dB to 30dB.
Figure. 4 shows the BER performance over the SNR from 15dB to 30dB and each sample is averaged by 100,000 independent channel realizations. The proposed method also outperforms other methods except for MMSE method. At 30dB SNR, the adaptive InterpolateNet achieves a BER of 4.8% while HDS InterpolateNet has a BER of 5.12%. The MMSE method has a minimum NMSE value of 2.7%. It should be noted that BER raise to 33.79% for optical receivers that directly demodulate the received signals. Replacing InterpolateNet with high-complexity neural networks should further improve both the NMSE and BER performance. This is because the neural network capacity is small due to the low complexity property of InterpolateNet and the deployed optical wireless channels are weakly frequency-selective type.
@@ V Conclusion
This paper proposes an effective and low-complexity neural network solution to robustly estimate optical wireless channels, which are possible to be frequency-selective. Compared with conventional neural networks, the proposed method selects one suitable candidate neural network based on the prediction of another neural network on the channel to provide the most accurate prediction. From the simulation results, the proposed solution maintains a coherent performance over randomly-generated optical channels and achieves superior performance on both NMSE and BER than other methods which have no prior channel information known. Compared to the HDS InterpolateNet and the LS estimate, the adaptive solution achieves reductions in NMSE and BER of 1.04% and 1.5% for the HDS InterpolateNet, and 11.71% and 20% for the LS estimate, respectively, at 20 dB SNR. Using the proposed solution does not require any actual channel information at the receiver end to precisely predict randomly-generated channels.

### v3

@@ IV Simulation Results
For the rest of this paper, $M$ is assumed to be 2 which represents a two-path channel for common indoor scenarios. Normalized mean square error (NMSE) is a key performance metric that evaluates the distance between the actual channel and the estimate of channel in the slot, which is defined as
$\mathrm{NMSE}(\mathbf{\hat{H}_{Slot}},\mathbf{H})=\frac{\mathbb{E}\left\{\left\Arrowvert\mathbf{\hat{H}_{\mathrm{Slot}}}-\mathbf{H}\right\Arrowvert_{F}^{2}\right\}}{{\mathbb{E}\{\left\Arrowvert\mathbf{H}\right\Arrowvert_{F}^{2}\}}},$ (10)
where $\left\Arrowvert\cdot\right\Arrowvert_{F}^{2}$ is the Frobenius norm. The bit error ratio (BER) is another performance metric.
TABLE I: System settings for optical wireless channel
@@
The offline training dataset for the three InterpolateNet networks (LDS, MDS or HDS channels) consists of 100,000 channel samples, 95% for training and 5% for validation. Referring to the environment parameters given in Table. I, we set $\mathbf{h_{\mathrm{LDS}}}$ = 1e-4 $\times$ [6.4, 0.21930, 0.09676, 0.06175, 0.04517] and $\mathbf{h_{\mathrm{HDS}}}$ = 1e-4 $\times$ [5.5, 0.30126, 0.13441, 0.08609, 0.06310] respectively for this indoor scenario. We use MSE loss for training InterpolateNets with Adam optimizer. The initial learning rate is 0.0002 and is reduced by 0.3 for every 10 epochs, for up to 100 epochs. The mini-batch size is 64 and L2 regularization item is $10^{-9}$.
@@ IV-A NMSE performance on randomly generated channels
We first evaluate the NMSE performance on randomly generated channels over both signal-to-noise ratio (SNR) and time, to investigate the robustness to allowed SNRs. The test channels are generated randomly according to Table. I.
@@
(b) NMSE performance over 90 seconds time (20dB SNR).
@@
Figure. 2a shows the NMSE performance over the SNR from 15dB to 30dB and each sample is averaged by 100,000 independent channel realizations. The proposed method is shown to have superior performance than other methods except for MMSE method. At 30dB SNR, the adaptive InterpolateNet achieves a NMSE of 0.5046 which is 0.0020 lower than that of HDS InterpolateNet. The MMSE method has a minimum NMSE value of 0.5001. For figure. 2b, the channel will alter among LDS, MDS and HDS for each 10 seconds and each sample is averaged over 100 independent channel realizations. Both the proposed method and the InterpolateNet trained with HDS condition outperform the LS estimate but are worse than MMSE method. The proposed method provides a better NMSE performance with a mean value of 0.5034, which is lower than 0.5052 for the InterpolateNet trained with HDS condition. Compared to the MMSE estimate with NMSE of 0.5012, the proposed solution does not require any actual channel information for online implementation resulting in poorer estimate performance and the NMSE variance of the proposed method is reasonably small over time.
@@ IV-B BER performance on randomly generated channels
We also evaluate the BER performance of each method on randomly generated channels over SNR.
Figure 3: BER results of LS, MMSE, direct signal detection, HDS-Net and the proposed method over SNR from 15dB to 30dB.
Figure. 3 shows the BER performance over the SNR from 15dB to 30dB and each sample is averaged by 100,000 independent channel realizations. The proposed method also outperforms other methods except for MMSE method. At 30dB SNR, the adaptive InterpolateNet achieves a BER of 4.8% while HDS InterpolateNet has a BER of 5.12%. The MMSE method has a minimum NMSE value of 2.7%. It should be noted that BER raises to 33.79% for optical receivers that directly demodulate the received signals. Replacing InterpolateNet with high-complexity neural networks should further improve both the NMSE and BER performance. This is because the neural network capacity is small due to the low complexity property of InterpolateNet and the deployed optical wireless channels are weakly frequency-selective type.
@@ V Conclusion
This paper proposes an adaptive low-complexity neural network solution to robustly estimate multi-path optical wireless channels. This proposed method selects one suitable candidate neural network based on the pre-estimate to provide an accurate prediction. From the simulation results, the proposed offline-trained solution performs robustly over randomly-generated optical channels and achieves better performance on both NMSE and BER than other methods which have no prior channel information known. Compared to the HDS-InterpolateNet and the LS estimate, the adaptive solution achieves reductions in NMSE and BER of 1.04% and 1.5% and 11.71% and 20% at 20 dB SNR respectively.